Chen Group acquired 20% of the outstanding ordinary shares of Cho Ltd. on December 31, 2019. The purchase price was ¥120000000 for 50000 shares. Cho declared and paid an ¥70 per share cash dividend on June 30 and on December 31, 2020. Cho reported net income of ¥74000000 for 2020. The fair value of Cho’s shares was ¥2600 per share at December 31, 2020.

At what amount is the investment reported on the statement of financial position under each of these methods at December 31, 2020? What is the effect on net income reported in 2020 under each of these methods?
Answer #1

Solution:

Reporting of investment on statment of financial position:

Fair value method =$ 120,000,000 +$3,000,000,000 =$3,120,000,000

Equity method = ($120,000,000 -3,500,000-3,500,000+14,800,000 ) =$127

,800,000

Reporting net income on the income staement:

Fair value method = Dividends revenue = (3,500,000 +3,500,000)=7,000,000

Equity method = Revenue from investment =14,800,000

working :

1)fair value / Urealized holding gain or loss calculation

Cost                                         = 120,000,000

Fair value (2600*120,000,000)=3,120,000,000

Unrealized gain or loss            =3,000,000,000

2)Equity investment (Dividend Revenue June 30)= 50,000 *$70 = 3,500,000

Equity investment (Dividend revenue December 31)= 50,000 *$70 = 3,500,000

revenue from investment = $74,000,000 *20%=$14,800,000